OpenStudy (anonymous):

Choose the equation of the line passing through the point (2, 6) and parallel to y = -3x - 4.

OpenStudy (anonymous):

the slope of this line will be -3, and the equation will be y-6 = -3(x-2) y=-3x+12

OpenStudy (anonymous):

Thank you so much! can you show me how to do it so I may finnish my test? if it's not trouble for you that is.

OpenStudy (anonymous):

ok. you know the slope of 2 parallel lines is the same. and the slope of the original line is -3 because its in the form y=mx+b where m is the slope. knowing this, you can use the point-slope formula: y-y1=m(x-x1) where (x1, y1) is a point the new line passes through. plug in m, x1, and y1 to get the equation for the new line

OpenStudy (anonymous):

oh, Thank you so much!